What is the percent ionization of a monoprotic weak acid solution that is 0.194 M? The acid-dissociation (or ionization) constant, Ka, of this acid is 4.92 × 10 − 10 .
Answer #1

[H+] =sqrt (Ka x C)

= sqrt (4.92 x 10^-10 x 0.194)

= 9.77 x 10^-6 M

percent ionization = H+ x 100 / C

= 9.77 x 10^-6 x 100 / 0.194

   = 0.00504%